# Measuring Fitness and Importance of Species in Food Webs

### v1

@@ Abstract
Ecosystems face intensifying threats from climate change, overexploitation, and other human pressures, emphasizing the urgent need to identify keystone species and vulnerable ones. While established network-based measures often rely on a single metric to quantify a species’ relevance, they overlook how organisms can be both carbon providers and consumers, thus playing a dual role in food webs. Here, we introduce a novel approach that assigns each species two complementary scores—an importance measure quantifying their centrality as carbon source and a fitness measure capturing their vulnerability. We show that species with high importance are more likely to trigger co-extinctions upon removal, while high-fitness species typically endure until later stages of collapse, in line with their broader prey ranges. On the other hand, low fitness species are the most vulnerable and susceptible to extinctions. Tested on multiple food webs, our method outperforms traditional degree-based analyses and competes effectively with eigenvector-based approaches, while also providing additional insights. Relying solely on interaction data, the approach is scalable and avoids reliance on expert-driven classifications, offering a cost-effective tool for prioritizing conservation efforts.
@@ I Introduction
In contemporary ecology, the preservation of ecosystems is an urgent and critical task. In this context, key challenges are understanding the repercussions of species loss and identifying the most vulnerable species. The significance of this task is further accentuated by the escalating anthropogenic threats, including climate change and over-exploitation of species, which strongly stress ecosystems [1, 2, 3, 4, 5, 6, 7, 8]. Species extinctions can trigger cascading effects throughout the ecosystem, generating a chain of secondary extinctions, thus amplifying the initial damage [9, 10]. To mitigate the effects of species removal, it is then crucial to quantify species “relevance” and “robustness”. This information can guide policies, focusing investments to preserve the most fundamental and less robust species.
@@
As in many complex systems, extinctions result from non-linear complex interactions and structures [11, 12], where each component is critical for the functioning of the ecosystem. Similar behaviors are also observed, for instance, when studying supply chains, where the removal of a single firm may result in a severe output loss of the entire economy [13, 14]. In this context, the network approach has emerged as a highly effective tool to study this type of structures and to comprehend the interplay between species or, more generally, any type of interacting entities. In this representation, each species is depicted as a node, while the links between nodes symbolize their interactions, such as mutualistic, host-parasite, or trophic relationships [15, 16]. Networks thus serve as mathematical representations that map the relationships within an ecological system, encompassing various types of connections. Among the different types of networks, food webs offer a representation of “who consumes whom” interactions and have been extensively studied from the network science perspective [17, 18, 19, 20, 21, 22, 23, 24]. Network measures, especially those related to PageRank or eigenvector centrality [25, 26, 27], have demonstrated substantial efficacy in quantifying the relevance of species in food webs [28, 29], identifying those whose removal would trigger the most significant impact on overall biodiversity. In particular, network measures prove particularly invaluable when considering large food webs, where exhaustive consideration of all potential extinction sequences is computationally prohibitive. However, while these approaches help detect crucial species, they fail to identify the most vulnerable ones. Indeed, many species play little or no role in the food web economy, but could still be in severe danger. This is, for instance, the case for pandas. Their extinction would not cause a direct secondary cascade in the food web, but their very non-diverse diet makes them extremely susceptible to changes in the ecosystem.
In light of these considerations, our research introduces a novel methodology for measuring a species’ dual nature in food webs: both as carbon source and consumer. Indeed most species absorb carbon from other species but are also consumed by larger predators or, eventually, by decomposers. This methodology quantifies the importance (susceptibility to predation and relevance for the food web) and fitness (predatory prowess and robustness to extinctions) of each species. By employing these two quantitative measures, we move beyond traditional one-dimensional evaluations that fail to capture the complex dual role of species in food webs. Conversely, our approach provides a holistic visualization of the entire food web, encapsulating its complex structure within a bidimensional plane, where each species is represented as a point. Our analysis demonstrates that important species are also those leading to larger extinction cascades, while low-fitness species possess little resistance to such events. Our methodology, therefore, enables a more accurate identification of both critical and vulnerable species within ecosystems, with the potential of greatly enhancing the effectiveness of conservation strategies.
@@
Figure 1: Cypress food web: network representations and Fitness-Importance plane. a, Directed monopartite subnetwork illustrating predator-prey relationships. b, Equivalent undirected bipartite representation of the subnetwork. c, Fitness-Importance plane depicting species’ ecological roles. Points represent individual species, with colors indicating functional categories and sizes proportional to species biomass. Selected species are labeled for reference. This visualization reveals the dual roles of species as both predators and prey within the ecosystem.
@@ II Results
II.1 The Fitness and Importance of Species
@@
Quantifying species’ roles and impacts within ecosystems presents a significant methodological challenge in ecological research. Traditional network-based methodologies approach this task by looking at the food web formed by the species. This consists of a directed network, where nodes are the species, while edges signify carbon exchange between these species. Being predation one of the main carbon exchange mechanisms, these graphs are often also called prey-predator networks. We show in Fig. 1a a schematized food web, where we adopt the convention of arrows going from predators to prey. Once a food web has been reconstructed, the standard approach consists in applying a centrality measure to rank the species. For instance, eigenvector centrality has proven very effective in determining which species have a stronger potential to cause chain extinctions if removed from an ecosystem [28]. These measures, however, are unable to assess which species are most in danger during these extinction cascades.
@@
To address this limitation, we introduce a novel algorithmic approach that quantifies species’ dual roles through two distinct measures: importance and fitness. Importance measures how important in the food web a species is, while fitness assesses a species’ predatory capabilities and adaptability within the food web. The importance of a species is high if it serves as prey for multiple species with low fitness scores. On the other hand, fitness evaluates a species’ predatory efficiency by assessing both the quantity and importance of its prey species. Species that prey upon diverse taxa, particularly those of low importance, demonstrate high predatory capacity and consequently receive high fitness values. This bi-dimensional characterization enables a more comprehensive understanding of species’ ecological roles compared to traditional network-based metrics.
Mathematically, we formulate this framework as a bi-dimensional non-linear map. Let $\bm{M}$ represent the adjacency matrix of the food web. The element $M_{ij}$ equals one if there is carbon transfer (predation) from species $i$ to $j$, and zero otherwise. The fitness ($F_{i}$) and importance ($I_{i}$) of species $i$ are computed through the following iterative map:
@@
$\left\{\begin{array}[]{lcl}F_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ji}\,/\,{I_{j}^{(n)}}\\ I_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ij}\,/\,{F_{j}^{(n)}}\end{array}\right.F_{i}^{(0)}=I_{i}^{(0)}=1\leavevmode\nobreak\ \forall i.$ (1)
@@
Here, $F_{i}^{(n)}$ and $I_{i}^{(n)}$ denote the fitness and importance of species $i$ at iteration $n$, respectively. The quantity $\delta$ is a regularization term that guarantees convergence. It does not affect the final ranking, as soon as it is way smaller than $\mathbf{M}$’s elements. In this paper, we set it at $10^{-3}$. The algorithm iterates until the convergence of both quantities. The first equation is telling us that a species has high fitness if it can absorb carbon by several sources, in particular from those with low importance. The latter are indeed only consumed by a small set of other species and are thus expected to be harder prey. The second equation implies that the importance of a species is high if it is consumed by several other species, in particular those with low fitness. Low fitness species can indeed only absorb carbon from a limited set of sources, so the extinction of any of those would put them in serious danger. For this reason, we identify vulnerability as the inverse of fitness. The lower the fitness of a species, the more vulnerable to food web shocks it is expected to be.
This approach shares notable similarities with the Economic Fitness and Complexity (EFC) algorithm developed for bipartite networks [30, 31, 32, 33]. While the EFC algorithm has successfully analyzed various bipartite nested networks, including ecological systems [16], its application to mono-partite directed networks represents a novel extension [34]. In ecological contexts, nestedness contributes significantly to ecosystem stability, biodiversity, and resilience, particularly in mutualistic networks where specialist species typically interact with generalist ones [35]. This similarity between the two approaches is not surprising, since, as shown in Fig. 1b, a food web can also be represented as a bipartite network. Differently from standard bipartite networks, layers contain the same set of species. The first looks at their role as predator, and the second one looks at their role as prey. Links between the two layers connect predators to their prey.
II.2 Fitness-Importance Plane
@@
Figure 2: Fitness-Importance plane. In royal blue, the marine ecosystems; in khaki, the terrestrial one; in medium aquamarine the mixed ones. Values are log-transformed and normalized to [0,1] for cross-food web comparison. The plane is divided into four quadrants. Base Trophic Nodes (top left quadrant), Carbon Bridge Nodes (top right quadrant), Sheltered Trophic Nodes (bottom left quadrant), Apex Trophic Nodes (bottom right quadrant).
@@
Our algorithm allows us to place all the species within an ecosystem on the 2-dimensional space defined by the fitness and the importance. Each axis captures a different aspect of species, namely their importance as prey and carbon suppliers and their prowess as predators and carbon absorbers. As an example, we show in Fig. 1c the fitness-importance plane for the Cypress dry season food web [36]. In this figure, each point corresponds to a species, different colors/shapes correspond to five different categories of species (see Methods for details), while point size is proportional to the biomass. Remarkably, species are not randomly scattered on the plane but tend to cluster following the classification. Producers and Organic Matter (dark green circles) are mostly placed on the top left corner since they are crucial carbon suppliers (high importance) and they do not absorb carbon from other species (low fitness). Conversely, large predators are placed in the bottom right corner, since they are both very hard to prey on and very good at preying. Also, a large fraction of secondary consumers are located in this region. The top right corner contains species that can absorb carbon from several other species, but that, at the same time, also transfer carbon to many other species. Not surprisingly in this area we find invertebrates and detritivores or mammals like rats. Finally, the bottom left corner mostly contains herbivores/omnivores and secondary consumers. These are species such as lizards that are not central in the food chain functioning, being hard to prey on, but with little to none preying capabilities.
Food web In-degree (%) Eigenvector-in (%) Importance (%)
@@
Table 1: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance rankings at the outset and keep them fixed throughout the simulated extinction process. Values represent means and standard deviation. For eigenvector-in and importance, the standard deviation is smaller than the last significative digit displayed. Higher percentages indicate greater ecosystem sensitivity to species loss based on the respective ranking method. Bold values indicate the best-performing method for each ecosystem.
As we show in Fig. 2, we can then approximately divide the fitness-importance plane into four regions:
  • Base Trophic Nodes (top left quadrant)
    These species are characterized by poor predatory ability but are heavily preyed upon by a multitude of species. They form the foundation for energy and nutrient flow, providing a primary source of food for consumers. They are both very important and very vulnerable (low fitness).
  • Carbon Bridge Nodes (top right quadrant)
    These species excel at absorbing or transforming carbon (often from detrital or non-living sources) and are widely consumed by other species. They “bridge” carbon from lower-level or non-living pools into higher trophic levels. Also them are very importance, but not as vulnerable, since their high fitness allows them to absorb carbon from very diverse sources.
  • Sheltered Trophic Nodes (bottom left quadrant)
    These species are neither strong predators nor frequent prey; they have defenses, behaviors, or ecological niches that protect them from heavy predation. They have low fitness, so high vulnerability, but their extinction would likely not cause cascades.
  • Apex Trophic Nodes (bottom right quadrant)
    These species possess high predatory ability with low vulnerability. They regulate prey populations and structure ecosystem dynamics from the top of the food web. They have low importance, but also low vulnerability.
Fig. 2 is realized considering six different ecosystems from terrestrial, aquatic, and mixed domains. Note that we normalized fitness and importance to compare species from different ecosystems (details in Methods). Colors denote the ecosystem type, while symbols show the different food webs. We observe a predominance of species in the Base Trophic Nodes ($42.09\%$) and Apex Trophic Nodes ($32.65\%$) quadrants, only a few Carbon Bridge Nodes ($2.30\%$) and a minority of Sheltered Tropic Nodes ($22.96\%$). We report in SI, for each ecosystem, the detailed list of all species in each quadrant. It is important to stress that fitness and importance only present a marginal anti-correlation, confirming that the two quantities actually capture complementary species’ properties.
II.3 Cascading Extinctions and Robustness of Species
Figure 3: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log fitness of extinct species vs. extinction step, based on importance ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log fitness vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison. This plot illustrates ecosystem resilience patterns and differential impacts of species loss across varied food web structures.
@@ II.3 Cascading Extinctions and Robustness of Species
The picture suggested by the fitness-importance plane, while reasonable, needs stronger validation. In particular, we need to understand whether the two measures we introduced are correlated with the predatory capabilities of animals and their vulnerability and importance as prey. These are not easy quantities to measure, but we can indirectly assess them.
@@
We start studying importance, which, as mentioned, is a measure of a species’ relevance as a carbon source. High-importance species play a major role in food webs since they represent the base of the food chain. This role can be quantified looking at the so called extinction curve of a food web. Such a curve is obtained by progressively removing species from the food web and looking at which other species get disconnected, thus remaining without any supplier. More details are reported in the Methods section. The premise is that a more accurate measure of species relevance within a food web should lead to a faster network collapse when species are removed in order of their ranking. Indeed, Eq. (1) implies that a species with high importance is preyed on by several low-fitness species. Therefore, its removal would likely cause many of these low-fitness, low-predatory capabilities species to be unable to sustain themselves. We quantify this effect using the mean extinction area (area under the extinction curve), where higher values indicate a better ability to identify critical species within the ecosystem. Table 1 presents the mean extinction areas following 3 different ranking strategies: in-degree, eigenvector-in centrality, and importance. We establish the initial rankings and keep them fixed throughout the simulated extinction process, whereas in the Supplementary Information, we present the scenario where rankings are recalculated after each node removal, including extinction curves plots for both cases across all food webs. The results reveal that both eigenvector-in and importance consistently outperform in-degree centrality across all studied ecosystems. Notably, eigenvector-in centrality and importance demonstrate comparable efficacy, with each measure excelling in different ecosystems. These results validate importance, confirming its ability to identify the most crucial species within food webs.
@@
As a second step, we study the ability of fitness to measure the vulnerability of species. A high-fitness species can prey on many other organisms, including those with low importance, i.e., those that are very hard to prey on. As a consequence, we expect high-fitness species to be less affected by extinctions within the ecosystem. They are indeed very diversified and less susceptible to such events. On the other hand, low-fitness species are expected to be the most vulnerable and thus the first to get extinct during cascade events. We can measure this by looking at the average (log) fitness of species that get extinct at a given step of the extinction curve construction. As shown in Fig. 3a for the cypress food web, this average increases with the step, meaning that high-fitness species only get extinct when most other species are already gone. We report next to each point a sample species going extinct in that step, while the inset shows the same plot, but using the average eigenvector-out centrality to measure (inverse) vulnerability. A much weaker correlation is observed. We repeat the same analysis for all the six food webs we considered, reporting the results in Fig. 3b. Note that we normalized both the step and the fitness for comparison reasons. In all cases, we observe the same behavior. More quantitatively, we can compute the Spearman correlation between the removal step and the (average log) fitness of the removed species. The average of this quantity among all ecosystems is $0.98$, while the average of the eigenvector-out centrality is $0.72$, thus consistently lower. We repeat the analysis using eigenvector-in centrality for species removal, yielding similar results that support fitness as an effective measure of species’ predatory skills and extinction resilience. In the Supplementary Information, we report details on Spearman correlation coefficients and eigenvector-in-based removals, along with additional comparisons between fitness, importance, and other centrality measures.

### v2

# Species Vulnerability and Ecosystem Fragility: A Dual Perspective in Food Webs

@@ Abstract
Ecosystems face intensifying threats from climate change, overexploitation, and other human pressures, emphasizing the urgent need to identify keystone species and vulnerable ones. While established network-based measures often rely on a single metric to quantify a species’ relevance, they overlook how organisms can be both carbon providers and consumers, thus playing a dual role in food webs. Here, we introduce a novel approach that assigns each species two complementary scores—an importance index quantifying their centrality as carbon source and a predatory index capturing their vulnerability. We show that species with high importance index are more likely to trigger co-extinctions upon removal, while high-robustness index species typically endure until later stages of collapse, in line with their broader prey ranges. On the other hand, low robustness index species are the most vulnerable and susceptible to extinctions. Tested on multiple food webs, our method outperforms traditional degree-based analyses and competes effectively with eigenvector-based approaches, while also providing additional insights. This scalable and data-driven approach, relying solely on interaction data, provides a cost-effective tool that complements expert classifications for prioritizing conservation efforts.
@@ I Introduction
The preservation of ecosystems is one of the most pressing challenges in contemporary ecology. A key aspect of this challenge lies in understanding the consequences of species loss and identifying the most vulnerable species. These tasks are increasingly urgent due to escalating anthropogenic pressures, such as climate change and over-exploitation, which severely disrupt ecosystems [1, 2, 3, 4, 5, 6, 7, 8]. Species extinctions often lead to cascading effects throughout the ecosystem, triggering secondary extinctions that amplify the initial damage [9, 10]. To mitigate these impacts, it is essential to quantify species’ ”relevance” and ”robustness,” guiding conservation policies to prioritize the preservation of fundamental and fragile species.
@@
Ecosystems, like many complex systems, are characterized by non-linear interactions and intricate structures [11, 12], where each component contributes to overall functionality. Similar dynamics are observed in other domains, such as interbank loan markets or supply chains, where the removal of a single bank or firm can significantly disrupt economic output [13, 14, 15]. In this context, network science has emerged as a powerful tool for studying such systems. By representing species as nodes and their interactions—such as mutualistic, host-parasite, or trophic relationships—as links, networks provide a mathematical framework for analyzing ecological systems [16, 17]. Among these networks, food webs, which depict ”who consumes whom” interactions, have been extensively studied to understand ecosystem dynamics [18, 19, 20, 21, 22, 23, 24, 25]. Network-based measures, particularly those derived from PageRank or eigenvector centrality [26, 27, 28], have proven effective in quantifying species’ relevance within food webs. These metrics identify species whose removal would have the most significant impact on biodiversity [29, 30]. This approach is especially useful for large food webs, where exhaustive simulations of all extinction scenarios are computationally infeasible. However, while these methods excel at identifying critical species, they often overlook those that are highly vulnerable. For instance, some species may play a minor role in the food web economy but remain extremely susceptible to extinction due to specialized diets or habitat requirements—pandas being a notable example.
While network science has provided numerous metrics for quantifying species’ roles in food webs, including measures of centrality and structural roles [31, 32], and the concept of keystone species has been explored through various network approaches [33], our methodology offers a distinct perspective by explicitly and simultaneously quantifying two complementary dimensions for each species: the importance index (measuring its centrality as a carbon source and its relevance within the food web) and the robustness index (capturing its predatory ability and resilience to extinction). Unlike single-index measures that characterize different aspects of species’ network roles [34, 35], this bi-dimensional framework provides a direct and separate assessment of these two fundamental ecological facets for each species, building upon the idea that considering both a species’ dependence on resources and its impact on its prey is crucial for a comprehensive understanding of its ecological role, as also explored through combinations of centrality indices [36]. This bi-dimensional characterization then allows us to visualize the food web’s structure, where each species is represented as a point on this plane. Our findings reveal that highly important species often drive significant extinction cascades when removed, whereas low- robustness index species exhibit limited resilience to such events. This methodology provides a more nuanced understanding of both critical and vulnerable species within ecosystems. By capturing the complex interplay between predation and survival in food webs, our approach offers a valuable, ecologically-informed perspective that could contribute to the broader framework of conservation strategies, potentially enhancing the prioritization of biodiversity protection efforts.
@@
Figure 1: Cypress food web: network representations and robustness index-importance index plane. a, Directed monopartite subnetwork illustrating predator-prey relationships. b, Equivalent undirected bipartite representation of the subnetwork. c, Robustness index-importance index plane depicting species’ ecological roles. Points represent individual species, with colors indicating functional categories and sizes proportional to species biomass. Selected species are labeled for reference. This visualization reveals the dual roles of species as both predators and prey within the ecosystem.
@@ II Results
II.1 The Robustness index and Importance index of Species
@@
Quantifying species’ roles and impacts within ecosystems presents a significant methodological challenge in ecological research. Traditional network-based methodologies approach this task by looking at the food web formed by the species. This consists of a directed network, where nodes are the species, while edges signify carbon flow between these species. As predation represents a major pathway for carbon flow between species, these networks are frequently termed prey-predator networks. We show in Fig. 1a a schematized food web, where we adopt the convention of arrows going from predators to prey. Once a food web has been reconstructed, the standard approach consists in applying a centrality measure to rank the species. For instance, eigenvector centrality has proven very effective in determining which species have a stronger potential to cause chain extinctions if removed from an ecosystem [29]. These measures, however, are unable to assess which species are most in danger during these extinction cascades and may not fully capture species’ vulnerability or the nuances of direct and indirect interactions (see also mixed trophic impact analysis [37] and loop analysis [38]).
@@
To address this limitation, we introduce a novel algorithmic approach that quantifies species’ dual roles through two distinct measures: the importance index and the robustness index. The importance index measures how important in the food web a species is, while the robustness index assesses a species’ predatory capabilities and adaptability within the food web. The importance index of a species is high if it serves as prey for multiple species with a low robustness index. On the other hand, the robustness index evaluates a species’ predatory efficiency by assessing both the quantity and the importance index of its prey species. Species that prey upon diverse taxa, particularly those of low importance index, demonstrate high predatory capacity and consequently receive high robustness index values. Mathematically, we formulate this framework as a bi-dimensional non-linear map. Let $\mathbf{M}$ represent the adjacency matrix of the food web. The element $M_{ij}$ equals one if there is carbon transfer (predation) from species $i$ to $j$, and zero otherwise. The robustness index ($R_{i}$) and importance index ($I_{i}$) of species $i$ are computed through the following iterative map:
@@
$\left\{\begin{array}[]{lcl}R_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ji}\,/\,{I_{j}^{(n)}}\\ I_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ij}\,/\,{R_{j}^{(n)}}\end{array}\right.R_{i}^{(0)}=I_{i}^{(0)}=1\leavevmode\nobreak\ \forall i.$ (1)
@@
Here, $R_{i}^{(n)}$ and $I_{i}^{(n)}$ denote the robustness index and the importance index of the species $i$ at iteration $n$, respectively. The quantity $\delta$ is a regularization term that guarantees convergence. It does not affect the final ranking, as soon as it is way smaller than $\mathbf{M}$’s elements. In this paper, we set it at $10^{-3}$. The algorithm iterates until the convergence of both quantities. The first equation tells us that a species has a high robustness index if it can absorb carbon from several sources, in particular from those with a low importance index. Referring to the second equation, species with a low importance index are consumed by few species, which suggests they are harder to prey upon. Conversely, the second equation implies that the importance index of a species is high if it is consumed by several other species, particularly those with a low robustness index. Low- robustness index species can indeed only absorb carbon from a limited set of sources, so the extinction of any of those would put them in serious danger. For this reason, we identify vulnerability as the inverse of the robustness index. The lower the robustness index of a species, the more vulnerable to food web shocks it is expected to be.
@@
Figure 2: Robustness index-importance index plane. In royal blue, the marine ecosystems; in khaki, the terrestrial one; in medium aquamarine the mixed ones. Values are log-transformed and normalized to [0,1] for cross-food web comparison. The plane is divided into four quadrants. Base Trophic Nodes (top left quadrant), Carbon Bridge Nodes (top right quadrant), Sheltered Trophic Nodes (bottom left quadrant), Apex Trophic Nodes (bottom right quadrant).
This approach shares notable similarities with the algorithms developed for bipartite networks in the field of Economic Complexity [39, 40, 41, 42], specifically with the algorithms of Economic Fitness and Complexity (EFC) [40, 41]. While the EFC algorithm has successfully analyzed various bipartite nested networks, including ecological systems [17], its application to mono-partite directed networks represents a novel extension [43]. In ecological contexts, nestedness contributes significantly to ecosystem stability, biodiversity, and resilience, particularly in mutualistic networks where specialist species typically interact with generalist ones [44]. This similarity between the two approaches is not surprising, since Fig. 1b illustrates how the directed trophic interactions of a food web can be visually represented using a bipartite layout, which can be useful for specific analytical approaches. Differently from standard bipartite networks, layers contain the same set of species. The first looks at their role as predator, and the second one looks at their role as prey. Links between the two layers connect predators to their prey.
II.2 Robustness index-Importance index Plane
@@
Our algorithm allows us to place all the species within an ecosystem on the 2-dimensional space defined by the robustness index and the importance index. Each axis captures a different aspect of species, namely their importance as prey and carbon suppliers and their prowess as predators and carbon absorbers. As an example, we show in Fig. 1c the robustness index-importance index plane for the Cypress dry season food web [45]. In this figure, each point corresponds to a species, different colors/shapes correspond to five different categories of species (see Methods for details), while point size is proportional to the biomass. Remarkably, species are not randomly scattered on the plane but tend to cluster following the classification. Producers and Organic Matter (dark green circles) are mostly placed on the top left corner since they are crucial carbon suppliers (high importance index) and they do not absorb carbon from other species (low robustness index). Conversely, large predators are placed in the bottom right corner, since they are both very hard to prey on and very good at preying. Also, a large fraction of secondary consumers are located in this region. The top right corner contains species that can absorb carbon from several other species, but that, at the same time, also transfer carbon to many other species. Not surprisingly in this area we find invertebrates and detritivores or mammals like rats. Finally, the bottom left corner mostly contains herbivores/omnivores and secondary consumers. These are species such as lizards that are not central in the food chain functioning, being hard to prey on, but with little to none preying capabilities.
In Fig. 2, the robustness index-importance index plane for each species is shown across all food webs. The plane is then divided into four regions:
  • Base Trophic Nodes (top left quadrant)
    These species are characterized by poor predatory ability but are heavily preyed upon by a multitude of species. They form the foundation for energy and nutrient flow, providing a primary source of food for consumers. They are both very important and very vulnerable (low robustness index).
  • Carbon Bridge Nodes (top right quadrant)
    These species excel at absorbing or transforming carbon (often from detrital or non-living sources) and are widely consumed by other species. They “bridge” carbon from lower-level or non-living pools into higher trophic levels. They are also very important, but not as vulnerable, since their high robustness index allows them to absorb carbon from very diverse sources.
  • Sheltered Trophic Nodes (bottom left quadrant)
    These species are neither strong predators nor frequent prey; they have defenses, behaviors, or ecological niches that protect them from heavy predation. They have a low robustness index, so high vulnerability, but their extinction would likely not cause cascades.
  • Apex Trophic Nodes (bottom right quadrant)
    These species possess high predatory ability with low vulnerability. They regulate prey populations and structure ecosystem dynamics from the top of the food web. They have a low importance index, but also low vulnerability.
Fig. 2 is realized considering six different ecosystems from terrestrial, aquatic, and mixed domains. Note that we normalized the robustness index and the importance index to compare species from different ecosystems (details in Methods). Colors denote the ecosystem type, while symbols show the different food webs. We observe a predominance of species in the Base Trophic Nodes ($42.09\%$) and Apex Trophic Nodes ($32.65\%$) quadrants, only a few Carbon Bridge Nodes ($2.30\%$) and a minority of Sheltered Tropic Nodes ($22.96\%$). We report in SI, for each ecosystem, the detailed list of all species in each quadrant. It is important to stress that the robustness index and importance index only present a marginal anti-correlation (average Spearman $s=-0.60\pm 0.18$ across the six food webs), confirming that the two quantities capture complementary species’ properties. This moderate anti-correlation is notably lower than the correlation we observe between eigenvector in-centrality and importance index (e.g., $s=0.93$ for the Cypress food web, as shown in Supplementary Information Section 3), indicating that the robustness index and importance index provide more distinct information.
II.3 Cascading Extinctions and Robustness of Species
Food web In-degree (%) Eigenvector-in (%) Importance index (%)
@@ II.3 Cascading Extinctions and Robustness of Species
Table 1: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance index rankings at the outset and keep them fixed throughout the simulated extinction process. Values represent means and standard deviation. For eigenvector-in and importance index, the standard deviation is smaller than the last significative digit displayed. Higher percentages indicate greater ecosystem sensitivity to species loss based on the respective ranking method. Bold values indicate the best-performing method for each ecosystem.
@@
The picture suggested by the robustness index-importance index plane, while reasonable, needs stronger validation. In particular, we need to understand whether the two measures we introduced are correlated with the predatory capabilities of animals and their vulnerability and importance as prey. These are not easy quantities to measure, but we can indirectly assess them.
@@
We start studying the importance index, which, as mentioned, is a measure of a species’ relevance as a carbon source. High-importance index species play a major role in food webs since they represent the base of the food chain. This role can be quantified looking at the so called extinction curve of a food web. Such a curve is obtained by progressively removing species from the food web and looking at which other species get disconnected, thus remaining without any supplier. More details are reported in the Methods section. The premise is that a more accurate measure of species relevance within a food web should lead to a faster network collapse when species are removed in order of their ranking. Indeed, Eq. (1) implies that a species with a high importance index is preyed upon by several species of low robustness index. Therefore, its removal would likely cause many of these species with low predatory capabilities to be unable to survive. We quantify this effect using the mean extinction area (area under the extinction curve), where higher values indicate a better ability to identify critical species within the ecosystem. Table 1 presents the mean extinction areas following 3 different ranking strategies: in-degree, eigenvector-in centrality, and importance index. We establish the initial rankings and keep them fixed throughout the simulated extinction process, whereas in the Supplementary Information, we present the scenario where rankings are recalculated after each node removal, including extinction curves plots for both cases across all food webs. The results reveal that both the eigenvector-in and importance index consistently outperform the degree-centrality across all ecosystems studied. Notably, eigenvector-in centrality and importance index demonstrate comparable efficacy, with each measure excelling in different ecosystems. These results validate the importance index, confirming its ability to identify the most crucial species within food webs.
Figure 3: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log robustness index of extinct species vs. extinction step, based on importance index ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log robustness index vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison. This plot illustrates ecosystem resilience patterns and differential impacts of species loss across varied food web structures.
@@
As a second step, we study the ability of the robustness index to measure species’ vulnerability. A high- robustness index species can prey on many other organisms, including those with a low importance index, i.e., those that are very hard to prey on. As a consequence, we expect high- robustness index species to be less affected by extinctions within the ecosystem. They are indeed very diversified and less susceptible to such events. On the other hand, low- robustness index species are expected to be the most vulnerable and, thus, the first to get extinct during cascade events. We can measure this by looking at the average (log) robustness index of species that get extinct at a given step of the extinction curve construction. As shown in Fig. 3a for the cypress food web, this average increases with the step, meaning that high- robustness index species only get extinct when most other species are already gone. We report next to each point a sample species going extinct in that step, while the inset shows the same plot, but using the average eigenvector-out centrality to measure (inverse) vulnerability. A much weaker correlation is observed. We repeat the same analysis for all the six food webs we considered, reporting the results in Fig. 3b. Note that we normalized both the step and the robustness index for comparison reasons. In all cases, we observe the same behavior. More quantitatively, we can compute the Spearman correlation between the removal step and the (average log) robustness index of the removed species. The average of this quantity among all ecosystems is $0.98$, while the average of the eigenvector-out centrality is $0.72$, thus consistently lower. We repeat the analysis using eigenvector-in centrality for species removal, yielding similar results that support the robustness index as an effective measure of species’ predatory skills and extinction resilience. In the Supplementary Information, we report details on Spearman correlation coefficients and eigenvector-in-based removals, along with additional comparisons between robustness index, importance index, and other centrality measures.